Operation Summary for January 10, 2011 - January 12, 2011 |
|
Monday Jan. 10 The day shift began with Operations monitoring the Tevatron (TeV) store 8413 (184.35E30) and the Recycler stash (97.3E10), with the Antiproton Source (Pbar) stacking (10.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported that the MI-60 upper bend power supply tripped off on a transformer temperature indication. Power supply technicians were able to get the supply back up within a half hour. Other short down times were caused by a MI RF vacuum trip, a Linac quadrupole power supply trip, MI gap clearing Kicker trips, and a Linac RF Buncher power supply trip. Operations disabled the T-978 beam at 1:30 PM when the experiment went into access. During the evening shift, Operations terminated store 8413 at 9:20 PM, put the TEV into HEP shot setup, and established store 8414 at 10:43 PM with an average initial luminosity of 326.0E30. The only recorded downtime was due to one minor Linac RF trip. Tuesday Jan. 11 The midnight shift began with Operations monitoring store 8414 (225.97E30) and the stash (71.6E10), with Pbar stacking (15.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that a MI RF station (MIRF15) tripped off on a modulator LCW flow fault that caused 15 minutes of downtime. A Linac Klystron RF station (KRF6) tripped off on a LCW flow fault. Operations established beam to T-978 at 7:30 AM. During the day shift, Operations terminated store 8414 at 2:36 PM, put the TEV into shot setup, and established store 8415 at 3:46 PM with an AIL of 313.9E30. Operations reported short downtimes caused by Linac quadrupole power supply trips and Linac RF Buncher trips. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations reported short downtimes caused by Linac quadrupole power supply trips and Linac RF Buncher trips. Wednesday Jan. 12 The midnight shift began with Operations monitoring store 8415 (117.82E30) and the stash (223.6E10), with Pbar stacking (13.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:25 AM. Operations terminated store 8415 at 7:24 AM and put the TEV into shot setup. Operations reported that NuMI tripped off a couple of times on a power supply ground fault; it reset okay. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 01/12/2011 email Fermilab |
| Security, Privacy, Legal | |